You find the following quotation for a stock:

Stock (ticker) Div Yld% P/E
(ttm)
High
(ttm)
Low
(ttm)
Volume
(100)
Last Change
AAPL 2.92 1.56 21.96 192.15 161.89 33,924 178.22 2.75

ttm stands for trailing twelve months.

Attempt 1/5 for 8 pts.

Part 1

What were earnings per share (EPS) over the most recent four quarters?
Answer #1

P/E = PRICE EARNINGS RATIO = MARKET PRICE/EPS

IN THE ABOVE TABLE, P/E = 21.96 AND MARKET PRICE (LAST TRADED) = 178.22

P/E = PRICE EARNINGS RATIO = MARKET PRICE/EPS

21.96 = 178.22/EPS

EPS = 178.22/21.96 = 8.1157 = 8.12 (ROUNDED TO 2 DECIMALS)

ANSWER : EPS = 8.12 (Thumbs up please)